You have just been hired by the EPA to help with the new gas mileage standards on new automobiles. The EPA would like to post average gas mileage as a 88% confidence interval. Suppose you collect data on 47 new Honda Accords and the average gas mileage for this sample was 28 with a sample standard deviation of 3.5. What is the HIGHER endpoint of your confidence interval? Be sure to use 2 decimal places.
Answer #1

Degrees of freedom = n - 1 = 47 - 1 = 46

Hence,

Critical t value for 88% confidence interval = 1.584

Therefore,

Higher endpoint = = 28.81